OpenStudy (anonymous):

Find the area of a square that is 0.21m on a side. (Round the area to the nearest hundredth)

OpenStudy (anonymous):

The area of a rectangle is always base * height, in this case, for a square the base and height are equal so, it is equal to the side^2

OpenStudy (anonymous):

i have these choices: 0.04m^2 0.42m^2 0.84.m^2

OpenStudy (mathstudent55):

A = s^2 A = (0.21 m)^2 A = 0.21 * 0.21 m^2

OpenStudy (anonymous):

so will it be 0.42m^2?

OpenStudy (mathstudent55):

Do you have a calculator? 0.21 * 0.21 is not the same as 0.21 + 0.21

OpenStudy (anonymous):

that would equal to 0.0441 if you multiply

OpenStudy (mathstudent55):

Correct. Since you are asked to round to the nearest 100th, the answer is 0.04 m^2

OpenStudy (anonymous):

thanks and can you help me with one more

OpenStudy (mathstudent55):

sure

OpenStudy (anonymous):

Subtracting Mixed numers 2 4/7-1 3/5

OpenStudy (mathstudent55):



OpenStudy (anonymous):

ok whats next

OpenStudy (mathstudent55):

You need a common denominator for the fraction parts. What is the least common denominator of 7 and 5?

OpenStudy (anonymous):

35?

OpenStudy (mathstudent55):

good. you need to change both fractions ot denominator 35